In 10 years A will be twice as old as B was 10 years ago. If A is now 9 years older than B then the present age of B is?
Answer by ikleyn(52781)

.
In 10 years A will be twice as old as B was 10 years ago. If A is now 9 years older than B then the present age of B is?
~~~~~~~~~~~~~~~~~~~~~~~~~

The condition says:

A + 10 = 2*(B-10),     (1)    ("In 10 years A will be twice as old as B was 10 years ago.")
A = B + 9.             (2)    (" A is now 9 years older than B:)

From (2), substitute the expression A = B + 9 into (1), replacing A. You will get

(B+9) + 10 = 2(B-10),

B + 19 = 2B - 20  --->  19 + 20 = B.

Hence, B is 39 years old now.

Then A = B + 9 = 39 + 9 = 48 years old now.

Solved.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".